OpenStudy (anonymous):

what does sin(beta/2) equal? 2sinbetacosbeta/2 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\sin\left(\frac{\beta}{2}\right)=\pm\sqrt{\frac{1-\cos(\beta)}{2}}\]